OpenStudy (anonymous):

How do I rewrite x^2 + 14x +49 in (a+b)^2 or (a-b)^2?

OpenStudy (anonymous):

(x+a)^2 = x^2 + a^2 + 2ax ex: (x+6)^2 = x^2 + 6^2 + 2*6x = x^2 + 36 + 12x

OpenStudy (anonymous):

Isnt it (x+7)^2 +14x?

OpenStudy (anonymous):

Very near :) Its (x+7)^2 (x+7)^2 = x^2 + 2*7x +7^2 = x^2 + 14x +49

OpenStudy (anonymous):

okay so for x^2-16x+64 it is (x+8)^2 -2*8x?

OpenStudy (anonymous):

If u see a negative sign, (x - a)^2 = x^2 + a^2 - 2ax ex: (x- 6)^2 = x^2 + 6^2 - 2*6x = x^2 + 36 - 12x x^2-16x+64 = (x - 8)^2

OpenStudy (anonymous):

That confused me... okay so you ignored the -16x in the last line

OpenStudy (anonymous):

First, you have to break it up into (x....)(x....) and then you must figure out the factors of 49 that make 14 when added. So you can make it into (x+7)(x+7). There is two of the same factors so you can make it into (x+7)^2

OpenStudy (anonymous):

lemme show it again: If u see a negative sign, (x - a)^2 = x^2 - 2ax + a^2 ex: (x- 6)^2 = x^2 - 2*6x+ 6^2 = x^2 - 12x + 36 (x - 8)^2 = x^2-2*8x+8^2 = x^2-16x+64

OpenStudy (anonymous):

Oh! okay thank you ! :)